The following reaction uses 3.42 g of nitrogen dioxide (NO2). What volume of dinitrogen tetroxide gas would be collected at 227 ̊C and 10.0 atm?
2 NO2 → N2O4
1
Expert's answer
2020-05-08T14:10:31-0400

"2NO_2 \rightarrow N_2O_4"


Convert g of "NO_2" to moles

"3.42 g (\frac{1 mol}{46.01g})=0.0743 mol"


Covert moles of "NO_2" to moles of "N_2O_4"

"0.0743 mol NO_2(\frac{1 mole N_2O_4}{2 moles NO_2})=0.0372 mol N_2O_4"



use Ideal Gas Law to find the volume of "N_2O_4"

"PV=nRT"

"V=\frac{nRT}{P}=\frac{0.0372\times0.082 \times (273+227)}{10.0}=0.153 L"